# Adapting Physics-Informed Neural Networks To Optimize ODEs in Mosquito Population Dynamics

### v1

@@ Abstract
Physics informed neural networks have been gaining popularity due to their unique ability to incorporate physics laws into data-driven models, ensuring that the predictions are not only consistent with empirical data but also align with domain-specific knowledge in the form of physics equations. The integration of physics principles enables the method to require less data while maintaining the robustness of deep learning in modeling complex dynamical systems. However, current PINN frameworks are not sufficiently mature for real-world ODE systems, especially those with extreme multi-scale behavior such as mosquito population dynamical modelling. In this research, we propose a PINN framework with several improvements for forward and inverse problems for ODE systems with a case study application in modelling the dynamics of mosquito populations. The framework tackles the gradient imbalance and stiff problems posed by mosquito ordinary differential equations. The method offers a simple but effective way to resolve the time causality issue in PINNs by gradually expanding the training time domain until it covers entire domain of interest. As part of a robust evaluation, we conduct experiments using simulated data to evaluate the effectiveness of the approach. Preliminary results indicate that physics-informed machine learning holds significant potential for advancing the study of ecological systems.
Keywords Physics-informed Neural Networks  $\cdot$ Hybrid Modelling  $\cdot$ Ecological Systems  $\cdot$ Ordinary Differential Equations  $\cdot$ Mosquito Population Modelling
Author Summary
Mosquitoes continue to rank among the world’s deadliest species. The ability to accurately simulate their seasonality and improve existing models is a crucial step towards timely risk assessment and subsequent planning of mitigation measures. Our solution extends existing research into the application of Physics-Informed Neural Networks for ordinary differential equation systems through the customization process that includes individual normalization and loss weight balancing. This has the effect of significantly improving training effectiveness and accuracy. This advanced training process includes temporal causality and optimized domain decomposition, facilitating practical applications, for example,in the field of biology. In a dual validation process, first using the Lorenz system and then, applying the system to mosquito population dynamics, our framework demonstrates robust applicability and offers valuable tools for complex ODE systems in various scientific fields.
1 Introduction
@@
1.1 Physics Informed Neural Networks
@@
1.2 Contribution and Paper Structure
While the majority of research on PINNs has concentrated on enhancing techniques for the effective training of Partial Differential Equations (PDEs), there has been less focus on customizing these methods for multi-variate, multi-equation Ordinary Differential Equation (ODE) systems. In this study, we explore the applicability and effectiveness of existing PINN methodologies to ODE systems. We introduce specific adjustments designed to optimize the training of ODE systems, including the individual normalization and loss weight balancing tailored for each variable and equation involved in the system. Moreover, we implement two additional steps in the training process to enhance both the effectiveness and accuracy of the models, and to simplify domain decomposition for practical application scenarios. We evaluate our modified framework using the Lorenz system, a classical model in dynamical systems theory, to demonstrate its applicability and value. Finally, we apply our approach to the modeling of mosquito population dynamics to validate its effectiveness in a practical biological context.
@@ 1.2 Contribution and Paper Structure
  • Our method includes a comprehensive normalization not only for inputs and outputs but also ODEs, ensuring more stable and accurate training processes in multi-variate multi-equation systems. In this respect, we implement an adaptive re-weighting of loss functions that individually adjusts the ODE loss weights, thus ensuring balanced training across physics constraints and data loss. Model training is further enhanced by incorporating a 3-phase progressive learning approach that respects temporal causality. This begins with data fitting for initialization, followed by cumulative training across sub-domains of increasing size, and informed neural networks guide mechanistic modeling from sparse experimental data finally tuning across the entire domain. In addition, we simplify domain decomposition, enforcing equality of boundary values at the interface between sub-domains and as a result, avoid extensive calculations to achieve continuity in high-order derivatives.
@@
The remainder of this paper is structure as follows: section 2 reviews the literature relevant to our study; section 3 introduces our PINN framework as applied to a system of ODEs and including our proposed extensions; in section 4, an ablation study using the Lorenz system is used to evaluate the effectiveness of each component in our framework; section 5 examines the wider impact and applicability of our approach by modeling mosquito populations; and finally in section 6, we conclude with some limitations of our current methods with suggestions for future work in this important research area.
2 Related Work
2.1 Mosquito Population Dynamics Modelling
@@
2.2 Normalization
Normalization is a crucial yet often overlooked step in the training of PINNs. In [53], researchers employed a strategy that involved dividing the input domains and applying individual input normalization alongside a unified global output normalization within their model computations. In [66], the authors suggested not only normalizing the inputs and outputs of the models but also non-dimensionalizing the differential equations integrated into the objective functions. In both [65, 34] and [64], both approaches added input- and output- scaling layers that multiply the inputs and outputs with their average magnitudes. However, this is performed at the model level and thus, still affects the objective functions and training efficiency. Despite recognizing the advantages of these approaches, there remains a lack of a systematic methodology for normalization, particularly in the context of dynamical systems characterized by multivariate system states and numerous differential equations. To address this in our research, we developed a normalization procedure that applies the MIN-MAX scheme to both the inputs and outputs of the neural networks, while appropriately transforming the system of ODEs. This procedure aids not only in aligning with the assumptions of model initialization but also in addressing the challenges associated with multi-scale and stiff issues commonly encountered in training PINNs.
2.3 Loss Re-weighting
@@ 2.4 Collocation points
We build on the methods presented in [47, 67] but crucially add two new steps: one for initial data fitting to improve model initialization (like [34, 64]) and include a final step which tunes across the entire domain.
2.5 Domain Decomposition
@@ 3 PINN Framework Development
3.1 PINN Structure
Consider $u(t)=\left(u^{(1)},u^{(2)},\dots,u^{(V)}\right)$ as a $V$-dimensional vector representing the state of a dynamical system at any given time $t$, where $t$ ranges from 0 to $T$. This dynamical system $u$ is governed by a set of $F$ ODEs as shown in equation 1, where …
@@ 3.1 PINN Structure
In the standard PINN framework [29], a neural network $U$ (parameterized by ${W_{U}}$), is used to approximate the solution $u$. This network, as a function defined in $[0,T]$, tries to estimate $u$ at any given time $t$, with $W_{U}$ being the trainable parameters of the network. In inverse problems where a few or all the parameters $\theta$ is not available, we can use a neural network, $\Theta^{(l)}$, to predict the unknown values $\theta_{l}$. For the sake of simplicity, we denote the set of all parameters, including the known or ones to be learnt by neural networks, as $\Theta$ and regard it as neural networks. The parameters $\Theta$, parameterized by $W_{\Theta}$, are also defined as functions in the domain $[0,T]$. If we let $W=\{W_{U},W_{\Theta}\}$, then the task now becomes an optimization problem when determining the parameters $W$ as shown in equation 2, which aims to minimize the multi-task objective function defined in equation 3.
@@
According to the universal approximation theorem [73], multi-layer perceptrons (MLP) [74] can approximate any continuous function on a given domain provided the MLP has sufficient complexity, as represented by the number of hidden layers and parameters in those hidden layers. It is shown that challenges in PINN may not be due to the capabilities of MLP [48]. For an MLP with $L$ layers, the mathematical formulation that describes its operation, layer by layer, is presented in equation 3.1 where $x$ is the input to the neural network, $h^{(l)}$ is the hidden state at layer $l$, $W^{(l)}$ and $b^{(l)}$ are the parameters of the layer $l$, and $\sigma$ is the activation function which provides non-linearity for the model.
@@
The "vanilla" PINN framework described above can approximate well when the ODE systems are relatively simple. However with more complex systems, especially ones that exhibit extreme stiffness, chaotic and multi-scale behavior, the basic setup tends to have difficulties in converging to a satisfactory local minimum. For this reason, we now proceed to describing our approaches to aid PINN training for both forward and inverse problem involved ODE systems.
3.2 ODE Normalization
Under normal circumstances, NNs output in the range [-1,1] and so for output values outside this range, data normalization plays a crucial role in the machine learning workflow, ensuring that both input and output variables remain within a reasonable range to enhance model convergence and accuracy. In the PINN context, this normalization process requires careful consideration, as any transformation results in modification to the corresponding ODE systems. In [34, 53], the authors consider normalization and de-normalization as parts of the overall model. However, this type of approach retains the original variable scales of the variables, potentially leading to significant imbalances in the objective function. [66] suggests normalizing the differential equations as well but its generalization is omitted. In this paper, we propose a systematic approach to do normalizations using PINNs involving ODEs. In particular, we utilize the MIN-MAX normalization scheme for both input and output variables, i.e. $t$, $u$ and $\theta$. To normalize the input variable, i.e. time $t$, we employ the transformation shown in equation 8, where $[T_{\min},T_{\max}]$ represents the time domain in which the models are trained. This will have the effect of normalizing the time variable $t$ to the range of $[-1,1]$.
@@ 3.2 ODE Normalization
To normalize the outputs from the neural networks, we define $\mathfrak{L}_{u},\mathfrak{U}_{u},\mathfrak{L}_{\theta}$, and $\mathfrak{U}_{\theta}$ as the lower and upper bounds for $u$ and $\theta$, respectively. When the difference between the lower and upper bounds is small, we adjust the bounds around the mean as $\mathfrak{L}=\min(\mathfrak{L},\mathfrak{M}-1)$ and $\mathfrak{U}=\min(\mathfrak{U},\mathfrak{M}+1)$ where $\mathfrak{M}=\frac{\mathfrak{L}+\mathfrak{U}}{2}$ is the middle point. It is important to note that these bounds are specific to each dimension in the dynamical system. These bounds can be estimated through collected data, inferred from domain knowledge, or estimated through approximate simulations when data is scarce. The normalization of the variables $u$ and $\theta$ at a normalized time $t^{\prime}$ is then obtained using equations 9 and 10 respectively.
@@
Following the normalization above, the transformed inputs and outputs, $t^{\prime}$, $u^{\prime}$ and $\theta^{\prime}$, now vary within the range [-1, 1]. Thus, instead of having neural networks approximating $u$ and $\theta$, it is instead more beneficial to use $U$ and $\Theta$ as surrogate models for $u^{\prime}$ and $\theta^{\prime}$. Following the normalization transformation in equations 9 and 10, it is necessary to adapt the loss functions accordingly. When computing $\frac{du^{\prime}}{dt}$ and applying 1, we arrive at the result shown in equation 11.
@@
Instead of using objective functions 4 and 5, we now consider the alternative equations 12 and 13 as these will provide re-scaling of the losses to the magnitude order of $u$ and $\frac{du}{dt}$, which are more similar to each other than the original objective functions are.
@@
3.3 Gradient Balancing
As this is a multi-task problem, we have different objective functions (e.g. Equation 3) as we are measuring different constraints. Previous research [36] highlighted that one point of failure in PINN training is the imbalance in gradients $\nabla_{W}\mathcal{L}_{data}$ and $\nabla_{W}\mathcal{L}_{f^{(i)}}$ in the update rule 6. It is observed that the differential equation residual loss dominates the overall loss due to the stiffness of the system. This makes the model prioritise optimizing the ODE constraint over matching the initial, boundary condition or data observations, leading the model to converge to a trivial solution, i.e. the null solution, violating conditions for the data. We address this issue by adopting a similar approach to [36] where balancing adjusts the weights $\lambda$ in the objective function based on statistics from the gradients $\nabla_{W}\mathcal{L}_{data}$ and $\nabla_{W}\mathcal{L}_{f^{(i)}}$. We extend this further to ODE systems by separately and individually assigning and adjusting the weights to each and every differential equation. In objective functions 3 and 6, the weights for data are set to 1 with component weights $\lambda_{i}$ each adjusted individually, to ensure that every equation in the system is given equal importance while aligning with the importance given to data. This extension to previous work is crucial because the scale and complexity of each equation are different and require different treatment during training.
@@ 3.3 Gradient Balancing
This gradient-balancing algorithm is described in Algorithm 1, where all weights $\lambda_{i}$ are initialized to 1 and are updated every $N$ steps. We compute the weight $\hat{\lambda}_{i}$ by calculating the ratio between the mean of absolute values of gradient $\nabla_{W}\mathcal{L}_{data}$ and the maximum of the absolute values for the gradients of $\nabla_{W}\mathcal{L}_{f^{(i)}}$. Due to the stochastic nature of gradient descent, the weight from these calculations can be highly volatile and thus, we update the weights $\lambda_{i}$ using the moving average formula in equation 15. The recommended values for hyper-parameters $\alpha$ and $N$ in the original study [36] are $\alpha=[0.5,0.9]$. However, we set $N=100$ and tune $\alpha$ in extreme cases $\alpha=0.99,0.9,0.5$, or set $N=1$ with $\alpha=0$.
3.4 Causal Training
@@
  • In the final tuning phase, we train the model using both loss terms, with the residual points from the entire time period. This phase is about refining and improving the solutions the model has learned so far, ensuring it follows the differential equations more accurately over the entire period.
The training process mainly happens during the second and third stages, with the first stage being the shortest. This is because neural networks can quickly learn to fit the given data too well. We usually limit the first phase, where the model learns to match the initial data, to between 5,000 and 10,000 steps. In the progressive causal phase, the model gradually learns to solve the differential equations over an expanding area. This phase takes longer because the model needs to successfully reduce its errors to a certain level before it can extend its learning to new areas. For systems that do not reliably converge, the number of steps for this phase should be sufficiently high. We set this phase to last between 50,000 and 100,000 steps to ensure thorough learning. Finally, in the fine tuning phase, we take the early stopping approach. We check the model’s performance every 1,000 steps with an evaluation loss, where all the weights $\lambda$s are set to 1.0. If no improvement is recorded in the model’s performance after several checks, the training ends. This ensures the model is as accurate as possible without wasting time on unnecessary training.
3.5 Domain Decomposition
@@
In particular, the approach divides the input domain into $S$ non-overlapped subdomains, $D_{s}=[T_{s-1},T_{s}],s=1,\dots,S$, with $T_{1}=0,T_{S}=T$. In each subdomain $D_{s}$, a neural network $U_{s}$ is defined, with the overall solution is defined in equation 16. With this definition, the goal is to train $U_{s}$ to approximate the solution $u$ in the subdomain $D_{s}$.
@@
Let $D_{s}^{*}=[T_{s-1}-O,T_{s}+O]$ be the extended subdomain of $D_{s}$, $O$ be 50% of overlapped size. We treat each subdomain $D_{s}^{*}$ as a separate PINN problem, where the model is trained using the framework described in section 3, including the normalization and gradient balancing measure. Starting from the first subdomain $D_{0}$, the model is trained with the initial condition data $\mathcal{D}_{u}$ provided by users. For subsequent subdomains $D^{*}_{s}$, the model is trained with data $\mathcal{D}_{u}$ generated by the previously trained model $U_{s-1}$ in the overlapped domain $[T_{s-1},T_{s-1}+O]$. The volume of data generated could be unlimited, we set the number from 10 to 100, depending on the granularity and the continuity across subdomains required. This process is repeated until the entire domain is covered. The final solution would be the combined predictions in equation 16.
@@
4 Ablation Study with Lorenz System
As different NN configurations or structures have been constructed for this task, we now evaluate the efficacy of the proposed PINN methodologies utilizing the Lorenz system, which consists of three coupled, nonlinear differential equations and is the common choice when evaluating ODEs [77]. This system is a classic example of chaotic behavior and is described in equation 20. Here, $x$, $y$, and $z$ represent the state variables of the system at any given time $t$ while $\sigma$, $\rho$, and $\beta$ are physical parameters of the system. We define two temporal domains for our experiments: a shorter duration with a maximum time of $T=2.0$ and a longer duration extending to $T=40.0$.
@@ 4 Ablation Study with Lorenz System
We conducted experiments on both forward and inverse problems using the Lorenz system, setting a shorter time frame of $T=2.0$ to test the first five models. The sixth model $\text{OdePINN}_{\text{grad+causal+domain}}$ is assessed on the forward problem over a longer duration with $T=20.0$.
Figure 1: Lorenz ODE system for the forward problem, with $U$ approximation of the system state. The blue line u_true represents the target for each of the 5 models used in the experiment. Graphs illustrate the performance across the x,y and z dimensions.
4.1 Forward Problem with T=2
@@
(a) Loss Terms across final models selecting through Early Stopping.
(b) ODE errors of the $\text{OdePINN}_{\text{grad}}$ model at different time $t$ during the training, motivating the need for Causal Training.
(c) The $x$-value Approximation Solution $U$ of the model $\text{OdePINN}_{\text{grad+causal}}$ at different steps during the training,
Figure 2: Loss analysis of OdePINN framework with Lorenz system, $t\in[0,2.0]$
Figure 1 presents the solution approximations produced by the five models, starting with the poorest performing models.

### v2

# Adapting Physics-Informed Neural Networks to Improve ODE Optimization in Mosquito Population Dynamics

@@ Abstract
Physics informed neural networks have been gaining popularity due to their unique ability to incorporate physics laws into data-driven models, ensuring that the predictions are not only consistent with empirical data but also align with domain-specific knowledge in the form of physics equations. The integration of physics principles enables the method to require less data while maintaining the robustness of deep learning in modelling complex dynamical systems. However, current PINN frameworks are not sufficiently mature for real-world ODE systems, especially those with extreme multi-scale behavior such as mosquito population dynamical modelling. In this research, we propose a PINN framework with several improvements for forward and inverse problems for ODE systems with a case study application in modelling the dynamics of mosquito populations. The framework tackles the gradient imbalance and stiff problems posed by mosquito ordinary differential equations. The method offers a simple but effective way to resolve the time causality issue in PINNs by gradually expanding the training time domain until it covers entire domain of interest. As part of a robust evaluation, we conduct experiments using simulated data to evaluate the effectiveness of the approach. Preliminary results indicate that physics-informed machine learning holds significant potential for advancing the study of ecological systems.
Introduction
@@
Physics Informed Neural Networks
@@
Contribution and Paper Structure
While the majority of research on PINNs has concentrated on enhancing techniques for the effective training of Partial Differential Equations (PDEs), there has been less focus on customizing these methods for multi-variate, multi-equation Ordinary Differential Equation (ODE) systems. In this study, we explore the applicability and effectiveness of existing PINN methodologies to ODE systems. We introduce specific adjustments designed to optimize the training of ODE systems, including the individual normalization and loss weight balancing tailored for each variable and equation involved in the system. Moreover, we implement two additional steps in the training process to enhance both the effectiveness and accuracy of the models, and to simplify domain decomposition for practical application scenarios. We evaluate our modified framework using the Lorenz system, a classical model in dynamical systems theory, to demonstrate its applicability and value. Finally, we apply our approach to the modelling of mosquito population dynamics to validate its effectiveness in a practical biological context.
@@
  • Our method includes a comprehensive normalization not only for inputs and outputs but also ODEs, ensuring more stable and accurate training processes in multi-variate multi-equation systems. In this respect, we implement an adaptive re-weighting of loss functions that individually adjusts the ODE loss weights, thus ensuring balanced training across physics constraints and data loss. Model training is further enhanced by incorporating a 3-phase progressive learning approach that respects temporal causality. This begins with data fitting for initialization, followed by cumulative training across sub-domains of increasing size, and informed neural networks guide mechanistic modelling from sparse experimental data finally tuning across the entire domain. In addition, we simplify domain decomposition, enforcing equality of boundary values at the interface between sub-domains and as a result, avoid extensive calculations to achieve continuity in high-order derivatives.
@@
The remainder of this paper is structured as follows: section Related Work reviews the literature relevant to our study; section Methods introduces our PINN framework as applied to a system of ODEs and including our proposed extensions; in section Ablation Study with Lorenz System, an ablation study using the Lorenz system is used to evaluate the effectiveness of each component in our framework; section Case Study-Based Validation examines the wider impact and applicability of our approach by modelling mosquito populations; and finally in section Discussion and Future Work, we conclude with some limitations of our current methods with suggestions for future work in this important research area.
Related Work
Mosquito Population Dynamics Modelling
@@
Normalization
Normalization is a crucial yet often overlooked step in the training of PINNs. In [53], researchers employed a strategy that involved dividing the input domains and applying individual input normalization alongside a unified global output normalization within their model computations. In [66], the authors suggested not only normalizing the inputs and outputs of the models but also non-dimensionalizing the differential equations integrated into the objective functions. In both [65, 34] and [64], both approaches added input- and output- scaling layers that multiply the inputs and outputs with their average magnitudes. However, this is carried out at the model level and thus, still affects the objective functions and training efficiency. Despite recognizing the advantages of these approaches, there remains a lack of a systematic methodology for normalization, particularly in the context of dynamical systems characterized by multivariate system states and numerous differential equations. To address this in our research, we developed a normalization procedure that applies the MIN-MAX scheme to both the inputs and outputs of the neural networks, while appropriately transforming the system of ODEs. This procedure aids not only in aligning with the assumptions of model initialization but also in addressing the challenges associated with multi-scale and stiff issues commonly encountered in training PINNs.
Loss Re-weighting
@@
We build on the methods presented in [47, 67] but crucially add two new steps: one for initial data fitting to improve model initialization (such as [34, 64]) and include a final step which tunes across the entire domain.
Domain Decomposition
@@ Methods
PINN Structure
Consider $u(t)=\left(u^{(1)},u^{(2)},\dots,u^{(V)}\right)$ as a $V$-dimensional vector representing the state of a dynamical system at any given time $t$, where $t$ ranges from 0 to $T$. This dynamical system $u$ is governed by a set of $F$ ODEs as shown in Eq (1), where …
@@
In the standard PINN framework [29], a neural network $U$ (parameterized by ${W_{U}}$), is used to approximate the solution $u$. This network, as a function defined in $[0,T]$, tries to estimate $u$ at any given time $t$, with $W_{U}$ being the trainable parameters of the network. In inverse problems where a few or all the parameters $\theta$ is not available, we can use a neural network, $\Theta^{(l)}$, to predict the unknown values $\theta_{l}$. For the sake of simplicity, we denote the set of all parameters, including the known or ones to be learnt by neural networks, as $\Theta$ and regard it as neural networks. The parameters $\Theta$, parameterized by $W_{\Theta}$, are also defined as functions in the domain $[0,T]$. If we let $W=\{W_{U},W_{\Theta}\}$, then the task now becomes an optimization problem when determining the parameters $W$ as shown in Eq (2), which aims to minimize the multi-task objective function defined in Eq (3).
@@
According to the universal approximation theorem [73], multi-layer perceptrons (MLP) [74] can approximate any continuous function on a given domain provided the MLP has sufficient complexity, as represented by the number of hidden layers and parameters in those hidden layers. It is shown that challenges in PINN may not be due to the capabilities of MLP [48]. For an MLP with $L$ layers, the mathematical formulation that describes its operation, layer by layer, is presented in Eq (PINN Structure) where $x$ is the input to the neural network, $h^{(l)}$ is the hidden state at layer $l$, $W^{(l)}$ and $b^{(l)}$ are the parameters of the layer $l$, and $\sigma$ is the activation function which provides non-linearity for the model.
@@
The ”vanilla” PINN framework described above can approximate well when the ODE systems are relatively simple. However with more complex systems, especially ones that exhibit extreme stiffness, chaotic and multi-scale behavior, the basic setup tends to have difficulties in converging to a satisfactory local minimum. For this reason, we now proceed to describing our approaches to aid PINN training for both forward and inverse problem involved ODE systems.
ODE Normalization
Under normal circumstances, NNs’ output is in the range [-1,1] and so for output values outside this range, data normalization plays a crucial role in the machine learning workflow, ensuring that both input and output variables remain within a reasonable range to enhance model convergence and accuracy. In the PINN context, this normalization process requires careful consideration, as any transformation results in modification to the corresponding ODE systems. In [34, 53], the authors consider normalization and de-normalization as parts of the overall model. However, this type of approach retains the original variable scales of the variables, potentially leading to significant imbalances in the objective function. [66] suggests normalizing the differential equations as well but its generalization is omitted. In this paper, we propose a systematic approach to do normalizations using PINNs involving ODEs. In particular, we utilize the MIN-MAX normalization scheme for both input and output variables, i.e. $t$, $u$ and $\theta$. To normalize the input variable, i.e. time $t$, we employ the transformation shown in Eq (8), where $[T_{\min},T_{\max}]$ represents the time domain in which the models are trained. This will have the effect of normalizing the time variable $t$ to the range of $[-1,1]$.
@@
To normalize the outputs from the neural networks, we define $\mathfrak{L}_{u},\mathfrak{U}_{u},\mathfrak{L}_{\theta}$, and $\mathfrak{U}_{\theta}$ as the lower and upper bounds for $u$ and $\theta$, respectively. When the difference between the lower and upper bounds is small, we adjust the bounds around the mean as $\mathfrak{L}=\min(\mathfrak{L},\mathfrak{M}-1)$ and $\mathfrak{U}=\min(\mathfrak{U},\mathfrak{M}+1)$ where $\mathfrak{M}=\frac{\mathfrak{L}+\mathfrak{U}}{2}$ is the middle point. It is important to note that these bounds are specific to each dimension in the dynamical system. These bounds can be estimated through collected data, inferred from domain knowledge, or estimated through approximate simulations when data is scarce. The normalization of the variables $u$ and $\theta$ at a normalized time $t^{\prime}$ is then obtained using Eqs (9) and (10) respectively.
@@
Following the normalization above, the transformed inputs and outputs, $t^{\prime}$, $u^{\prime}$ and $\theta^{\prime}$, now vary within the range [-1, 1]. Thus, instead of having neural networks approximating $u$ and $\theta$, it is instead more beneficial to use $U$ and $\Theta$ as surrogate models for $u^{\prime}$ and $\theta^{\prime}$. Following the normalization transformation in Eqs (9) and (10), it is necessary to adapt the loss functions accordingly. When computing $\frac{du^{\prime}}{dt}$ and applying Eq (1), we arrive at the result shown in Eq (11).
@@
Instead of using objective functions (4) and (5), we now consider the alternative Eqs (12) and (13) as these will provide re-scaling of the losses to the magnitude order of $u$ and $\frac{du}{dt}$, which are more similar to each other than the original objective functions are.
@@
Gradient Balancing
As this is a multi-task problem, we have different objective functions (e.g. Eq (3)) as we are measuring different constraints. Previous research [36] highlighted that one point of failure in PINN training is the imbalance in gradients $\nabla_{W}\mathcal{L}_{data}$ and $\nabla_{W}\mathcal{L}_{f^{(i)}}$ in the update rule (6). It is observed that the differential equation residual loss dominates the overall loss due to the stiffness of the system. This makes the model prioritise optimizing the ODE constraint over matching the initial, boundary condition or data observations, leading the model to converge to a trivial solution, i.e. the null solution, violating conditions for the data. We address this issue by adopting a similar approach to [36] where balancing adjusts the weights $\lambda$ in the objective function based on statistics from the gradients $\nabla_{W}\mathcal{L}_{data}$ and $\nabla_{W}\mathcal{L}_{f^{(i)}}$. We extend this further to ODE systems by separately and individually assigning and adjusting the weights to each and every differential equation. In objective functions (3) and (6), the weights for data are set to 1 with component weights $\lambda_{i}$ each adjusted individually, to ensure that every equation in the system is given equal importance while aligning with the importance given to data. This extension to previous work is crucial because the scale and complexity of each equation are different and require different treatment during training.
@@
This gradient-balancing algorithm is described in Algorithm 1, where all weights $\lambda_{i}$ are initialized to 1 and are updated every $N$ steps. We compute the weight $\hat{\lambda}_{i}$ by calculating the ratio between the mean of absolute values of gradient $\nabla_{W}\mathcal{L}_{data}$ and the maximum of the absolute values for the gradients of $\nabla_{W}\mathcal{L}_{f^{(i)}}$. Due to the stochastic nature of gradient descent, the weight from these calculations can be highly volatile and thus, we update the weights $\lambda_{i}$ using the moving average formula in Eq (15). The recommended values for hyper-parameters $\alpha$ and $N$ in the original study [36] are $\alpha=[0.5,0.9]$. However, we set $N=100$ and tune $\alpha$ in extreme cases $\alpha=0.99,0.9,0.5$, or set $N=1$ with $\alpha=0$.
Causal Training
@@
  • In the final tuning phase, we train the model using both data loss and ODE losses, with the ODE residual points drawn at random with uniform distribution across the entire domain. The goal here is to refine and improve the solutions the model has learned so far, ensuring it follows the differential equations more accurately over the entire period.
The training process mainly happens during the second and third stages, with the first stage being the shortest. This is because neural networks can quickly learn to fit the given data too well. We usually limit the first phase, where the model learns to match the initial data, to between 5,000 and 10,000 steps. In the progressive causal phase, the model gradually learns to solve the differential equations over an expanding area. This phase takes longer because the model must successfully reduce its errors to a required level before it can extend its learning to new areas. For systems that do not reliably converge, the number of steps for this phase should be sufficiently high. We set this phase to last between 50,000 and 100,000 steps to ensure thorough learning. Then, in the final tuning phase, the early stopping approach is adopted. The model’s performance is checked every 1,000 steps with an evaluation loss, where all weights $\lambda$s are set to 1.0. If no improvement is recorded in the model’s performance after several checks, the training ends. This ensures the model is as accurate as possible without time spent on unnecessary training.
Domain Decomposition
@@
In particular, the approach divides the input domain into $S$ non-overlapped subdomains, $D_{s}=[T_{s-1},T_{s}],s=1,\dots,S$, with $T_{1}=0,T_{S}=T$. In each subdomain $D_{s}$, a neural network $U_{s}$ is defined, with the overall solution is defined in Eq (16). With this definition, the goal is to train $U_{s}$ to approximate the solution $u$ in the subdomain $D_{s}$.
@@
Let $D_{s}^{*}=[T_{s-1}-O,T_{s}+O]$ be the extended subdomain of $D_{s}$, $O$ be 50% of overlapped size. We treat each subdomain $D_{s}^{*}$ as a separate PINN problem, where the model is trained using the framework described in section Methods, including the normalization and gradient balancing measure. Starting from the first subdomain $D_{0}$, the model is trained with the initial condition data $\mathcal{D}_{u}$ provided by users. For subsequent subdomains $D^{*}_{s}$, the model is trained with data $\mathcal{D}_{u}$ generated by the previously trained model $U_{s-1}$ in the overlapped domain $[T_{s-1},T_{s-1}+O]$. The volume of data generated could be unlimited, we set the number from 10 to 100, depending on the granularity and the continuity across subdomains required. This process is repeated until the entire domain is covered. The final solution would be the combined predictions in Eq (16).
@@
Ablation Study with Lorenz System
As different NN configurations or structures have been constructed for this task, we now evaluate the efficacy of the proposed PINN methodologies utilizing the Lorenz system, which consists of three coupled, nonlinear differential equations and is the common choice when evaluating ODEs [77]. This system is a classic example of chaotic behavior and is described in Eq (20). Here, $x$, $y$, and $z$ represent the state variables of the system at any given time $t$ while $\sigma$, $\rho$, and $\beta$ are physical parameters of the system. We define two temporal domains for our experiments: a shorter duration with a maximum time of $T=2.0$ and a longer duration extending to $T=40.0$.
@@
Where a bespoke technique is not applied, conventional methods are used by default. Specifically, in the absence of ODE normalization, input and output variables retain their original scale; without Gradient Balancing, the weights $\lambda_{i}$ are set to a fixed value of 1; without Causal Training, collocation points are uniformly sampled across the domain; and without Domain Decomposition, the model is trained on the entire domain of interest in a single training. In this approach, experiments were conducted on both forward and inverse problems using the Lorenz system, setting a shorter time frame of $T=2.0$ to test the first five models. The sixth model $\text{OdePINN}_{\text{grad+causal+domain}}$ is assessed on the forward problem over a longer duration with $T=20.0$.
Forward Problem with T=2
@@
Fig 1 presents the solution approximations produced by the five models, starting with the poorest performing models.
@@
Fig 1: Lorenz ODE system for the forward problem, with $U$ approximation of the system state. The blue line u_true represents the target for each of the 5 models used in the experiment. Graphs illustrate the performance across the x,y and z dimensions.